A marketing firm spent $12,400 in advertisements in January for Alexa's used car emporium. February sales at Alex's were $341,000. In August, the firm spent $8,600 on advertisements, and sales in the following month were $265,000.
a. write a linear equation relating the amount spent on advertisement x with sales in the following month y. write your answer in slope- intercept form.
b. use the equation found in part a to predict the sales amount if 10,000 is spent on advertisements in the previous month.

Answer by mathtech5(4)

Set up the points: (12,400 , 341,000) and (86,000 , 265,000)
Find the slope: (y2 - y1)/ (x2 - x1)
(265000-341000) / (86000-12400) = 20
Use the equation y=mx+b: y = 20x + b
Put a point in for x and y: 265000 = 20(8600) + b
Now solve for b: 265000=172000 + b
Subtract 172000 from both sides to get: 93000 = b
The equation would be : y = 20x + 93,000
To predict the sales when 10,000 is spent on advertisements, substitute 10,000 in for x: y = 20(10000) + 93000 = 293000
The predicted sales would be: $293,000